Jack is twice as old as Lacy. In 3 years the sum of their ages will be 54. How old are they now?
Answer by stanbon(75887)

Jack is twice as old as Lacy.
J = 2L
---------
In 3 years the sum of their ages will be 54.
(J+3)+(L+3) = 54
------------------------
How old are they now?
-----
Substitute for J and solve for "L":
(2L+3)+(L+3) = 54
3L + 6 = 54
3L = 48
L = 16 (Lacy's age now)
---
Since J=2L, J = 32 (Jack's age now)
=========================================
Cheers,
Stan H.